OpenStudy (anonymous):

If y varies inversely as x, find the constant of variation if y = 3 when x = -2.

OpenStudy (anonymous):

\[y =\frac{ k }{ x },where k is constant.\] when x=-2,y=3 \[3=\frac{ k }{ -2 },k=-2\times3=-6,

OpenStudy (anonymous):

\[3=\frac{ k }{ -2},k=-2\times3=-6\]